Julie is laying on the ground and spots a bird up in the tree. The angle of elevation of Julie (on the ground) to the bird is 65◦. The bird is at the top of the 24 ft tree. To the nearest tenth of a foot, how far is Julie from the base of the tree?
Answer by ewatrrr(24785)

tan 65° = opp/adj = 24/x
x = 24/tan(65°)